Aт 8.0kg point mass and a 15.0kg point mass are held in place 50.0cm apart. A particle of mass m is released from a point between the masses, 20cm from the 8.0kg mass along a line connecting the two fixed masses. Find the magnitude and direction of the acceleration of the particle.


1
Expert's answer
2019-10-07T11:04:29-0400

Let us use notation:

"m_1 = 8 kg, m_2 = 15 kg" - masses of the particle 1 and 2 respectively, "m" - the mass of the middle particle.

"r = 50 cm" - distance between the particles,

"r_1 = 20 cm, r_2 = r - r_1 = 30 cm" - distances from the middle particle to the first and second respectively.


The first and second particles attract the middle particle with the force "F_{1,2} = G \frac{m_{1,2} m}{r_{1,2}^2}", according to Newton's law of universal gravity. The absolute value of the net force is "F = |F_1 - F_2| = G m \left| \frac{m_1}{r_1^2} - \frac{m_2}{r_2^2}\right|" . The value inside the modulus is approximately "33.3 \frac{kg}{m^2}", i.e. it is positive, so the net force is directed towards the first particle. The absolute value of acceleration is "a = \frac{F}{m} = G \left| \frac{m_1}{r_1^2} - \frac{m_2}{r_2^2}\right| \approx 2.22 \cdot 10^{-9} \frac{m}{s^2}" .